Problem 5-1 Simple Interest versus Compound Interest [LO1] First City Bank pays 8 percent simple interest on its savings account balances, whereas Second City Bank pays 8 percent interest compounded annually. If you made a $68,000 deposit in each bank, how much more money would you earn from your Second City Bank account at the end of 8 years? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Difference in accounts
Answer #1

Total amount at First City Bank after 8 years = 68,000 + 68,000*8%*8 = $111,520

Total amount at Second City Bank = 68,000(1.08)^8 = $125,863.25

More money at Second Bank = $125,863.25 - 111,520

= $14,343.25